OpenStudy (anonymous):

What values for θ(0≤θ≤2π) satisfy the equation 3sinθ = sinθ-1?

OpenStudy (anonymous):

Do some algebra on the equation first (simplify).... subtract sin(theta) from both sides, then divide by sin(theta) on both sides... what you got left as a result?

OpenStudy (anonymous):

I'm so confused.

OpenStudy (anonymous):

start with the original equation then subtract sin(theta) from both sides first... 3sinθ = sinθ-1 3sinθ - sinθ = sinθ-1- sinθ 2sinθ = -1 this is the first part. now divide both sides by sin(theta)...

OpenStudy (anonymous):

sorry... divide both sides by "2"....

OpenStudy (anonymous):

So sinθ = -1/2

OpenStudy (anonymous):

ok... do you know what angle(or angles) gives a -1/2 when you take the sine of that angle?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

have you used the unit circle?

OpenStudy (anonymous):

I have not.